# Artificial Intelligence HCI Human-Computer Interaction

### v2

@@ 4.1.3. Pains and Gains
Collaboration
Most participants reported that while collaboration and iterative discussions are essential to their QDA process, those can be tedious and frustrating, particularly in cases where multiple interpretations are possible or different backgrounds are present, making it difficult to reach consensus. A few participants described onboarding others and initial discussions as cumbersome.
On the positive side, participants highlighted the value of sharing their thoughts with others and thereby learning to better understand the data. These discussions were seen as enriching and helping to refine ideas.
@@
Participants with limited QDA experience often expressed concerns about the lack of structure and guidance in QDA. They reported feeling uncertain about whether they were conducting the analysis correctly, expressing a need for assistance, such as having a second person review their work. This was linked to the perception that coding is subjective and inherently messy, making it difficult to evaluate whether their observations had objective value or were merely personal insights.
@@
Our study provides a detailed view of how HCI researchers perform and perceive QDA in their everyday practice. Our interviews revealed that while participants’ workflows are similar in structure, they vary highly depending on the project, methodology, data type, degrees of cooperation, and approaches in which codebooks are created. As Figure 1 illustrates, deductive coding is a somewhat straightforward process, whereas inductive coding, more data-driven and interpretative, involves several loops. This is also mirrored in the pain points participants highlighted and has implications for the integration of AI into the QDA workflow, as highly unique use cases also need to be addressed. Similarly, other perceived “pains and gains” during the workflow are often not mutually exclusive: what one researcher considers as enjoyable and vital to the research can be another researcher’s biggest annoyance. For example, less experienced researchers expressed a need for guidance and structure in their workflows, reflecting the importance of mentorship and learning in QDA. This highlights the importance of building AI assistants in a configurable and customizable manner, not as a one-size-fits-all solution, but as systems that can adapt to researchers’ varying needs and levels of expertise.
@@ 4.2.4. Conditions and Requirements
Participants were generally open to AI in QDA, provided it enhanced rather than replaced their skills. They valued AI for offloading repetitive tasks, enabling them to focus on deeper interpretive work. Others saw an opportunity in AI to enrich the analysis with new views and perspectives. Hence, AI was perceived both as a tool for efficiency and as a catalyst for analysis. Participants expressed several concerns and mentioned conditions that had to be met. They suggested developing AI tools that operate offline to allow for greater control over the data’s use and storage. Quality and trustworthiness of AI outputs were another major concern, especially given AI’s current limitations in dealing with nuanced data and the risk of generating false information. Participants stressed the importance of rigorously testing and validating AI tools before integrating them into their research workflows to ensure their reliability in research contexts. Still, they saw potential in AI tools that are developed in alignment with their needs, with a focus on transparency, control, and ethical compliance. Notably, willingness to use AI did not correlate with participants’ experience level, suggesting these conditions are shared across different levels of expertise.
@@ 5. A Framework for AI-based Support that Aligns with HCI Researchers’ Real-Life QDA Workflows
Our results indicate a general openness to exploring the potential of AI support for QDA among HCI researchers. However, this willingness is balanced by significant concerns about research ethics, the quality of AI output, and maintaining control over the research process. Based on our results, we identify different roles AI could play within QDA from recurring themes from our interviews regarding the appropriate use of AI in QDA, researchers’ conditions and concerns, and the stages in the QDA workflow where AI could provide benefits. We contextualize these roles in related work. To guide the development of future AI tools and help researchers decide when and how to integrate AI into their workflows, we propose a framework outlining three levels of AI involvement: (a) minimal AI involvement (b) moderate AI involvement (c) high AI involvement We provide concrete examples of how AI can be integrated at various stages of the QDA workflow, summarized in Figure 2.
Figure 2. Our framework for integrating AI into QDA, depicting 3 stages from less to more AI involvement, visualizing possible roles for AI we deduced from the data for each stage.:
@@ 5.1. Minimal AI Involvement: Supporting Human-Centric Workflows
At this level, AI has little to no direct impact on the core QDA workflow and primarily serves as a supportive tool.
@@
At the lowest level, the researcher conducts the entire QDA process independently, without AI assistance. This is particularly relevant when data is highly sensitive, context-dependent, or when there’s a lack of trust in AI’s ability to achieve a nuanced understanding. While one participant was reluctant to use AI altogether, others were open to involving AI. Some researchers express skepticism about the use of AI for QDA, expressing concerns about opacity in AI processes (Khalili, 2023), ethical challenges (Ashwin et al., 2023), and the erosion of human interpretive insight (Hitch, 2024).
@@
AI is utilized to perform repetitive, non-interpretive tasks that do not require human judgment, allowing researchers to maintain full control over the interpretive aspects of QDA. As P12 noted, “transcription can be done by AI as long as it complies with our privacy practices”. AI can handle tasks like adjusting data formats, cleaning data, and organizing it for analysis. For instance, tools like Whisper [^5] https://openai.com/index/whisper/ can transcribe audio recordings.
@@
AI acts as a virtual assistant within QDA software, improving usability by guiding users through software features and tools. P1 explained, “some tools are difficult to get to know”. Examples of intelligent user interfaces include virtual assistants like Microsoft’s Clippy, providing extensive domain knowledge and maintaining a model of the user.
AI as an Intelligent User Interface [Uncaptioned image] Figure 3. AI provides step-by-step guidance within QDA software, suggesting functionalities based on user behavior and answering questions to support with problems in tool usage. :
@@
In this role, AI enhances productivity by automating evaluative tasks, such as summarizing data, quantifying patterns, or performing basic statistical analyses. P2 highlighted that “calculating IRR in MAXQDA can be a big pain. AI could offer a customizable IRR calculation”, enabling quicker insights. Beyond QDA, AI tools have demonstrated utility across diverse data work practices, including writing, translation, and literature review. Tools such as DeepL [^6] https://www.deepl.com/translator assist with seamless translation, while platforms like Rayyan [^7] https://www.rayyan.ai/ streamline systematic literature review processes, enabling researchers to focus on higher-level analytical tasks.
AI as a Productivity Tool [Uncaptioned image] Figure 4. AI generates summary reports of QDA projects, such as identifying the most or least frequent codes, or calculating IRR, allowing the researcher to review these insights and make informed decisions more efficiently. :
@@
AI can be used to train aspiring qualitative researchers through interactive, hands-on learning and assist in onboarding by familiarizing new team members with coding workflows and providing personalized training exercises. Our findings suggest that less experienced researchers, who often expressed uncertainty about their coding accuracy, could benefit from AI training tools that provide structured guidance and immediate feedback, addressing gaps in expertise. Palea et al. (Palea et al., 2024) propose a tool that generates AI hints to help students improve their coding by understanding annotation errors.
AI as a Training Tool [Uncaptioned image] Figure 5. AI supports new researchers in learning and practicing QDA through personalized, interactive instructions, real-time feedback, and simulated collaboration exercises to develop critical thinking and improve coding consistency. :
@@ 5.2. Moderate AI Involvement: Human-AI Collaboration
AI mediates between multiple coders, helping to reach consensus and facilitate collaboration. Complaining about time-consuming and tiring meetings with collaborators, P4 noted that AI could assist by focusing discussions on relevant parts and highlighting interpretive discrepancies, reducing the time and effort spent in meetings. For instance, Aemonium (Drouhard et al., 2017) highlights ambiguities and inconsistencies between multiple coders by providing visual overviews of assigned codes. CodeWizard (Ganji et al., 2018) identifies coders’ certainty and disagreements to improve consensus. CollabCoder (Gao et al., 2024) provides real-time synchronization, acting as a suggestion provider, mediator, and facilitator throughout the workflow.
AI as a Mediator [Uncaptioned image] Figure 6. AI helps researchers resolve coding conflicts and reach consensus efficiently by analyzing discrepancies and suggesting resolutions. :
@@
Participants often felt uncertain when conducting QDA and saw potential in using AI for validating their coding, e. g., by comparing their results with AI-identified codes or segments. In this role, AI serves as a validation tool, identifying gaps or inconsistencies, and providing a different perspective. By providing a layer of validation, such a tool can address researchers’ concerns on uncertainty and subjectivity in the coding process. For instance, AI could verify the completeness of coding and highlight passages that may need revisiting. Our findings indicate that this role is particularly valued by less experienced researchers, who expressed a greater need for external validation to build confidence in their analysis, emphasizing the potential for AI tools to support skill development and ensure reliability.
AI for Validation [Uncaptioned image] Figure 7. AI helps identify errors and inconsistencies while keeping researchers actively engaged, e.g., by suggesting potentially missing annotations and providing them for review. :
@@
AI supports refining the codebook by providing names, descriptions, and restructuring. P1 noted that AI could help organize the codebook once the initial data analysis is completed. P2 added that AI might suggest merging or grouping similar codes to increase clarity. For instance, MAXQDA’s AI features [^8] https://www.maxqda.com/blogpost/ai-coding-of-qualitative-data provide explanations for coded segments, aiding researchers in organizing and refining their codebooks.
AI for Refinement & Explanations [Uncaptioned image] Figure 8. AI provides support by offering concise code definitions and names, e.g., for the final codebook, and suiting quotes for codes. :
@@
AI provides data-driven insights to support the human-led coding process, e. g., by giving warnings or identifying patterns. For instance, Annota (Palea et al., 2024) marks passages for potential additional improvements.
AI for Analytics [Uncaptioned image] Figure 9. AI identifies patterns in coding and provides warnings to support decision-making, e.g., when codes overlap or themes could be grouped. :
@@
AI suggests codes based on its analysis, which researchers can accept, modify, or reject, maintaining control over the process. Our findings indicate that participants were generally open to receiving AI recommendations for new themes and suggestions on where to apply a code. This openness was particularly evident among participants working with large datasets or repetitive coding tasks. These suggestions keep the researcher in control while allowing AI to enhance the efficiency and breadth of the coding process. For example, Marathe et al. (Marathe and Toyama, 2018) and Rietz et al. (Rietz and Maedche, 2021) developed tools that provide coding suggestions based on simple rules, supporting the extension of the coding to unseen data. Gao et al. (Gao et al., 2024) explored collaborative workflows where LLMs propose new codes and code groups.
AI Suggestions [Uncaptioned image] Figure 10. AI analyzes the data and recommends codes which can be reviewed by the researcher and refined by the AI, preserving human agency. :
@@
AI and human researchers work together throughout the QDA process. AI assists at various stages, such as generating and refining codes, while the human researcher guides the overall process. Participants preferred an interactive or conversation-like AI, with P7 suggesting that AI could act as a “sparring partner” for discussing categories and collaboratively assigning codes. For instance, Chew et al. (Chew et al., 2023) propose a holistic role for LLMs in co-developing and revising codebooks and conducting final coding.
AI as a Sparring Partner [Uncaptioned image] Figure 11. Researchers and AI collaborate interactively throughout the QDA workflow, from codebook creation, over codebook refinement, to codebook application. :

### v3

@@ 4.1.3. Pains and Gains
Participants with limited QDA experience often expressed concerns about the lack of structure and guidance in QDA. They reported feeling uncertain about whether they were conducting the analysis correctly, expressing a need for assistance, such as having a second person review their work. This was linked to the perception that coding is subjective and inherently messy, making it difficult to evaluate whether their observations had objective value or were merely personal insights. Here, collaboration with other, more experienced researchers was brought up as helpful.
@@
Collaboration Across Stages
Most participants reported that while collaboration and iterative discussions are essential to their QDA process, those can be tedious and frustrating, particularly in cases where multiple interpretations are possible or different backgrounds are present, making it difficult to reach consensus. Participants emphasized that different analysts may apply categories based on different underlying assumptions (“understand categories differently,” P15), and that negotiating these differences can make the codebook “explode” rather than converge on shared definitions (P7). In several workflows, this alignment work involved iterative loops of independent coding followed by comparison meetings and codebook revisions; in some cases, this was driven by expectations around IRR, described as requiring very frequent coordination (e.g., meeting “nearly every day to get some IRR,” P2). Where IRR was calculated, participants reported that it could be discouraging; ranging from “only 50%” agreement that revealed diverging interpretations of the same code (P8) to experiences of extremely low agreement (“cruel 20–30%,” P7).
A few participants described onboarding others and initial discussions as cumbersome. This included the effort of briefing and calibrating how codes should be applied (P2) as well as the broader challenge that new coders may lack full project context (P1). Some teams addressed disagreement by involving a third person as a tie-breaker (P10; P12). More generally, dependency risks were also reported: relying on another person can mean delays if they drop out or do not deliver the expected quality (P12).
Several accounts also stressed that tooling can amplify friction: collaboration was described as difficult when researchers cannot work synchronously in the same project (P1; P7) or when sharing/merging is unstable or “a disaster” (P5), leading teams to rely on workarounds such as Google Sheets/Word memos to coordinate and document decisions (P1; P11; P4).
On the positive side, participants highlighted the value of sharing their thoughts with others and thereby learning to better understand the data and the QDA process. These discussions were seen as enriching and helping to refine ideas. Some explicitly described the interpretive exchange as enjoyable (“the discussion makes [it] fun,” P15; “I enjoy the process of discussing because that gives you another perspective,” P6), and valued collaboration as a second pair of eyes to improve clarity and consistency (P1; P9).
@@
Our study provides a detailed view of how HCI researchers perform and perceive QDA in their everyday practice. Our interviews revealed that while participants’ workflows are similar in structure, they vary highly depending on the project, methodology, data type, approaches in which codebooks are created, and degrees and quality of collaboration. As Figure 1 illustrates, deductive coding is a somewhat straightforward process, whereas inductive coding, more data-driven and interpretative, involves several loops. This is also mirrored in the pain points participants highlighted and has implications for the integration of AI into the QDA workflow, as highly unique use cases also need to be addressed. Similarly, other perceived “pains and gains” during the workflow are often not mutually exclusive: what one researcher considers as enjoyable and vital to the research can be another researcher’s biggest annoyance. For example, some participants described iterative discussions to reach consensus as tedious, while others experienced them as educational. For example, less experienced researchers expressed a need for guidance and structure in their workflows, reflecting the importance of mentorship and learning in QDA facilitated by collaboration with others. This highlights the importance of building AI assistants in a configurable and customizable manner, not as a one-size-fits-all solution, but as systems that can adapt to researchers’ varying needs and levels of expertise.
@@ 4.2.4. Conditions and Requirements
Participants were generally open to AI in QDA, provided it enhanced rather than replaced their skills. They valued AI for offloading repetitive tasks, enabling them to focus on deeper interpretive work. Others saw an opportunity in AI to enrich the analysis with new views and perspectives. Hence, AI was perceived both as a tool for efficiency and as a catalyst for analysis. Participants expressed several concerns and mentioned conditions that had to be met. They suggested developing AI tools that operate offline to allow for greater control over the data’s use and storage. Quality and trustworthiness of AI outputs were another major concern, especially given AI’s current limitations in dealing with nuanced data and the risk of generating false information. Participants stressed the importance of rigorously testing and validating AI tools before integrating them into their research workflows to ensure their reliability in research contexts. Still, they saw potential in AI tools that are developed in alignment with their needs, with a focus on transparency, control, and ethical compliance. Generally, they preferred AI to support collaborative sensemaking, whether in engagement with the data or other researchers, rather than replacing it. Notably, willingness to use AI did not correlate with participants’ experience level, suggesting these conditions are shared across different levels of expertise.
@@ 5. A Framework for AI-based Support that Aligns with HCI Researchers’ Real-Life QDA Workflows
Our results indicate a general openness to exploring the potential of AI support for QDA among HCI researchers. However, this willingness is balanced by significant concerns about research ethics, the quality of AI output, and maintaining control over the research process. Participants have expressed different suggestions and ideas for incorporating AI in their workflows. Our interviews revealed that researchers do not conceptualize AI support in QDA as a binary choice between automation and human interpretation. Instead, participants articulated a spectrum of acceptable AI involvement, ranging from complete exclusion to conditional or near-autonomous delegation. Based on these insights, we propose a tiered framework that characterizes AI integration along increasing levels of agency, epistemic influence, and autonomy. The framework is developed inductively from participants’ suggestions and grounded in existing tools and prior work to provide tangible design examples. Importantly, the tiers are not intended as a linear progression or normative recommendation. Instead, they represent a design space from which researchers may selectively adopt AI support depending on context, data sensitivity, methodological commitments, and personal trust. We group the identified roles into three levels of AI involvement – minimal, moderate, and high (see Table 2).
Table 2. A tiered framework for integrating AI into Qualitative Data Analysis (QDA), based on participant interviews, with three stages ranging from minimal to high AI involvement.
AI Role Origin Related Work Participant Quote
Minimal Human-only P3 (Davison et al., 2024) “Naturally, I want to do most of the things myself”
AI for basic technical tasks P2, P4, P6, P7, P12 (WhisperAI, [n. d.]) “I am very happy about automated transcriptions”
AI as an intelligent user interface P1 (Wikipedia contributors, 2025) “Some tools are really difficult to get to know.”
AI as a productivity tool P2 (DeepL SE, [n. d.]; Rayyan Systems Inc., [n. d.]) “Calculating IRR in MAXQDA can be a big pain.”
AI as a training tool P4, P1 (Palea et al., 2024; SocioCultural Research Consultants, LLC, [n. d.]; Wang et al., 2019) “Labeling exercises”
Moderate AI as a mediator P4 (Drouhard et al., 2017; Gao et al., 2024; Ganji et al., 2018) “Comparison of different codebooks”
AI for validation P12, P4 (Morgan, 2023) “You could use it to validate your own ideas – and still discard it.”
AI for refinement and explanations P1, P2, P8 (VERBI Software, [n. d.]) “AI can help to get some structure into my codebook”
AI for analytics P12, P5, P10, P9, P4 (Palea et al., 2024) “Give me warnings if I am too detail-oriented.”
AI suggestions P12, P8, P11, P4 (Marathe and Toyama, 2018; Rietz and Maedche, 2021; Gao et al., 2024) “The tool suggests an important code or a frequent word.”
AI as a sparring partner P7, P10, P11, P9 (Chew et al., 2023) “As a sparring partner, I can imagine this in all steps”
High AI with human-in-the-loop P8, P4 (Cohn et al., 2024; Dai et al., 2023) “The AI suggests a codebook, and I give feedback”
AI with human approval P13, P12, P8 (Drápal et al., 2023; Xiao et al., 2023; Bano et al., 2023) “AI could apply my codebook, and I review a random subsample”
AI with conditional autonomy P4 (Lennon et al., 2021) “I see AI as tooling for low-hanging fruit, so I can focus on more complex codes”
Full AI delegation (partial or full) P1, P8 (Lam et al., 2024; Rasheed et al., 2024; OpenAI, 2024) “The tool could automatically create a codebook.”
@@ 5.1. Minimal AI Involvement: Supporting Human-Centric Workflows
At this level, AI has little to no direct impact on the core QDA workflow and primarily serves as a supportive tool. We emphasize this tier as an important baseline rather than a limitation, as it highlights that AI-supported QDA must remain optional and that full human control continues to be a legitimate and valued practice.
@@
At the lowest level, the researcher conducts the entire QDA process without AI assistance. This approach is particularly relevant when working with sensitive data, highly contextual material, or when researchers lack trust in AI’s ability to achieve nuanced interpretation or simply do not have access to suitable tools. While one participant was reluctant to use AI altogether (P3), others were open to involving AI. Given longstanding concerns about opacity, ethical risks, and the erosion of interpretive depth in AI-supported qualitative work (Khalili, 2023; Ashwin et al., 2023; Hitch, 2024), this level remains crucial.
@@
More commonly, participants accepted AI for repetitive, non-interpretive tasks such as transcription, formatting, or data cleaning. As P12 noted, “transcription can be done by AI as long as it complies with our privacy practices”. For instance, tools like Whisper (WhisperAI, [n. d.]) can transcribe audio recordings.
@@
AI may also act as a usability layer within QDA software, guiding researchers through complex features without shaping analytic outcomes. P1 explained, “some tools are difficult to get to know” and likened this role to metaphors such as Microsoft’s Clippy (Wikipedia contributors, 2025), providing domain knowledge and maintaining a model of the user.
AI as an Intelligent User Interface [Uncaptioned image] Figure 2. AI provides step-by-step guidance within QDA software, suggesting functionalities based on user behavior and answering questions to support with problems in tool usage. Figure 3 visualizes Role 3, AI as an Intelligent User Interface. A pictogram of a person with a speech bubble asks: How can I merge two versions of my project? A pictogram of a computer with a speech bubble answers: Let me assist you in combining multiple versions of your project. First, drop both files in the document browser on the left. […].
@@
In this role, AI enhances productivity by automating evaluative tasks, such as summarization, translation, literature screening, or inter-rater reliability (IRR) calculations. P2 highlighted that “calculating IRR in MAXQDA can be a big pain. AI could offer a customizable IRR calculation”, suggesting AI could reduce overhead. Tools like DeepL(DeepL SE, [n. d.]) and Rayyan(Rayyan Systems Inc., [n. d.]) illustrate how AI already demonstrates its utility beyond QDA.
AI as a Productivity Tool [Uncaptioned image] Figure 3. AI generates summary reports of QDA projects, such as identifying the most or least frequent codes, or calculating IRR, allowing the researcher to review these insights and make informed decisions more efficiently. Figure 4 visualizes Role 4, AI as a Productivity Tool. A pictogram shows 2 paper sheets, labeled Human-coded data input. An arrow leads to a symbol of a computer, labeled AI summarization tool. From here, 3 arrows lead further. Arrow 1 leads to IRR. Arrow 2 leads to bar graph pictogram. Arrow 3 leads to a sum pictogram, labeled Reports.
@@
Finally, AI can support onboarding and training, particularly for less experienced qualitative researchers. P4 mentioned the idea of “labeling exercises”, guided practice to build confidence and consistency. For example, the Dedoose [^5] https://www.dedoose.com/userguide/interraterreliability training center supports research teams in maintaining coding consistency. Palea et al. (Palea et al., 2024) propose a tool that generates AI hints to help students improve their coding by understanding annotation errors. Wang et al. show AI can be used as teaching tool in data science (Wang et al., 2019).
AI as a Training Tool [Uncaptioned image] Figure 4. AI supports new researchers in learning and practicing QDA through personalized, interactive instructions, real-time feedback, and simulated collaboration exercises to develop critical thinking and improve coding consistency. Figure 5 visualizes Role 5, AI as Training Tools. 3 visualizations are separated by lines, depicting different examples. The first is titled Introduction and Learning. A speech bubble asks: How to start coding? A pictogram of a book is labeled as AI Tutor. The second is titled Guided Practice with Feedback. A pictogram of a paper sheet is framed by 3 pictograms of tags on the left, and 2 checkmarks and 1 X on the right. It is labeled Interactive Coding. The third is titled Simulated Collaboration and Review. A cloud-shaped bubble contains 3 pictograms of persons with paper sheets. A big paper sheet pictogram is depicted at the bottom. It is labeled AI as virtual team member.
@@ 5.2. Moderate AI Involvement: Human-AI Collaboration
Collaboration was envisioned by participants as being supported by AI by surfacing disagreements, aligning diverging codebooks, and focusing discussions during consensus finding. Rather than resolving meaning, AI highlights where interpretation diverges, reducing coordination costs. For instance, Aemonium (Drouhard et al., 2017) highlights ambiguities and inconsistencies between multiple coders by providing visual overviews of assigned codes. CodeWizard (Ganji et al., 2018) identifies coders’ certainty and disagreements to improve consensus. CollabCoder (Gao et al., 2024) provides real-time synchronization, acting as a suggestion provider, mediator, and facilitator throughout the workflow. Participants also imagined AI support for collaboration as a glitch-free shared workspace with reliable real-time synchronization, scaffolding for onboarding and early calibration of new team members, and as an additional collaborator or “third instance” to spar with (P4; P6). The latter included producing tentative codes to compare against human coding (P7). Still, they expressed skepticism that AI could replace theme-building collaborative conversations (P3) and questioned whose labels an AI should learn from when research teams already show low agreement (P12).
AI as a Mediator [Uncaptioned image] Figure 5. AI helps researchers resolve coding conflicts and reach consensus efficiently by analyzing discrepancies and suggesting resolutions. Figure 6 visualizes Role 6, AI as a Mediator. It depicts a small flowchart with 3 steps. Step 1 is labeled Researchers perform initial coding. It contains 3 pictograms of persons, labeled A, B, and C, each with a pictogram of a paper sheet. From each person, an arrow leads to step 2, a pictogram of a computer, labeled AI mediator analyzes coding discrepancies … . An arrow leads to step 3: A pictogram of a paper sheet framed with letters A, B, and C. It is labeled … and suggests possible resolutions. From here, an arrow leads back to step 1. The arrow is labeled Researchers discuss and agree on final codes.
@@
Participants, especially those less experienced, expressed uncertainty about their coding decisions saw potential in using AI for validation, e. g., by comparing their results with AI-identified codes or segments. In this role, AI serves as a validation tool, identifying gaps or inconsistencies. For instance, AI could verify the completeness of coding and highlight passages that may need revisiting (P4).
AI for Validation [Uncaptioned image] Figure 6. AI helps identify errors and inconsistencies while keeping researchers actively engaged, e.g., by suggesting potentially missing annotations and providing them for review. Figure 7 visualizes Role 7, Ai for Validation. It depicts a small flowchart with 4 elements. The start point is labeled Human coder completes initial task. It shows a pictogram of a person with a pictogram of a paper sheet, that is labeled Coded data segments. From here, an arrow leads to 2 elements representing different possible scenarios. Element 1 contains a pictogram of a computer, a pictogram of a codebook, and a pictogram of a person with a paper sheet. It is labeled AI scans for potentially missing annotations. Element 2 contains a symbol for a computer, a pictogram of 3 paper sheets labeled as Already coded data set, and a pictogram of person with a paper sheet. It is labeled AI checks for consistency across dataset.
@@
Once researchers have developed an initial understanding of the data, AI may assist with refining codebooks by suggesting clearer names, definitions, examples, or restructuring. For example, P2 noted that AI might suggest merging or grouping similar codes to increase clarity. For instance, MAXQDA’s AI features [^6] https://www.maxqda.com/blogpost/ai-coding-of-qualitative-data provide explanations for coded segments, aiding researchers in organizing and refining their codebooks.
AI for Refinement & Explanations [Uncaptioned image] Figure 7. AI provides support by offering concise code definitions and names, e.g., for the final codebook, and suiting quotes for codes. Figure 8 visualizes Role 8, AI for Refinement and Explanations. It depicts pictogram of a person, and a pictogram of a computer, conversing with each other. The conversation has 3 elements with speech bubbles for each. In the first element, the person asks: Given the following codebook and dataset, provide descriptions and examples for each code. The speech bubble depicts a book and a dataset. The computer answers with a codetable, consisting of 2 rows for Code A, Code B, and 2 columns titled Definition and Example. In the second element, the person asks: Given the following codebook and database, suggest shorter, clearer names for codes. The speech bubble depicts a book and a dataset. The computer answers with Code A, and an arrow leading to 3 dots to represent alternatives. In the third element, the person asks: Suggest codes that support the rational of this code. The speech bubble depicts a book and a dataset. The computer answers: Quote for A: ‘…’.
@@
Participants also described AI providing reflexive, data-driven insights, such as warnings about overly granular codebooks, expected reliability issues, or frequency patterns. P12 suggested an AI tool warning them about the codebook getting too long. P9 suggested a table that shows quantitative statistics on the frequency and patterns of each code, which can be used for further analysis, without taking it away from them. Taken together, these suggestions help researchers in a reflexive way where instead of generating reasons and justifications, they can still come up with their own reasons. For instance, Annota (Palea et al., 2024) marks passages for potential additional improvements.
AI for Analytics [Uncaptioned image] Figure 8. AI identifies patterns in coding and provides warnings to support decision-making, e.g., when codes overlap or themes could be grouped. Figure 9 visualizes Role 9, AI for Analytics. It depicts a small flowchart with 2 steps. Step 1 depicts a Codebook and annotated data, represented by pictograms of a paper sheet and a cylinder. An arrow leads to 3 speech bubbles. The first bubble says: These two codes might overlap: …. It is labeled Warnings. The second bubble says: These themes could be grouped: … . It is labeled Insights. The third bubble says: You changed a code. Review these passages: … . It is labeled Markers.
@@
AI suggests codes based on its analysis, which researchers can accept, modify, or reject, maintaining control over the process. Our findings indicate that participants were generally open to receiving AI recommendations for new themes and suggestions on where to apply a code. P8 ideated that “the tool suggests an important code, or a frequently mentioned word.” This openness was particularly evident among participants working with large datasets or repetitive coding tasks. Existing tools exemplify this role (Marathe and Toyama, 2018; Rietz and Maedche, 2021; Gao et al., 2024).
AI Suggestions [Uncaptioned image] Figure 9. AI analyzes the data and recommends codes which can be reviewed by the researcher and refined by the AI, preserving human agency. Figure 10 visualizes Role 10, AI Suggestions. It depicts a small flowchart with 3 elements. Element 1 is labeled AI suggests. It depicts a computer with a light bulb. 3 arrows point to element 2, a paper sheet with lines and 3 squares representing highlights. Element 3 is labeled Researcher reviews. It depicts a pictogram of a person, and a question mark, a checkmark, and an x next to the paper sheet. From here, an arrow leads back to element 1. It is labeled AI refines.
@@
Some described AI as a “sparring partner”, a conversational collaborator that proposes alternatives, challenges assumptions, and supports sensemaking throughout the workflow. Participants even described this as supporting the existing human collaboration as a third instance or sparring partner (P7; P6; P4), with AI sometimes mediating consensus work (also see AI as a Mediator). Notably, participants preferred an interactive or conversation-like AI (P7). For instance, Chew et al. (Chew et al., 2023) propose a holistic role for LLMs in co-developing and revising codebooks and conducting final coding.
AI as a Sparring Partner [Uncaptioned image] Figure 10. Researchers and AI collaborate interactively throughout the QDA workflow, from codebook creation, over codebook refinement, to codebook application. Figure 11 visualizes Role 11, AI as a Sparring Partner. It contains 3 elements with a progress arrow underneath. The 3 elements each depict 2 pictograms: a person, a computer, and a double arrow between them. The first element is labeled Codebook Creation. The second element is labeled Codebook Refinement. The third element is labeled Codebook Application.